The following equations refer to the goods market of an economy in billions of euros:

C=480 +0,5Yd

I=110

T=70

G=250

1.Solve for goods market equilibrium 

2.Find equilibrium disposable income

3.find equilibrium consumption 

4.calculate the private savings,public savings and investment spending


1
Expert's answer
2021-05-31T15:39:04-0400

Solution:

1.). Goods market equilibrium: S = I

First derive the saving function from the consumption function:

C = 480 + 0.5Yd = 480 + 0.5(Y – T)

S = -a + (1 – b) Y

S = -480 + (1 – 0.5) Y

I = 110

S = I

-480 + (1 – 0.5) Y = 110

-480 + 0.5Y = 110

0.5Y = 110 + 480

0.5Y = 590

Y = 590/0.5 = 1.180

Goods market equilibrium = 1,180

 

2.). At equilibrium disposable income: Y = AE

 

Y = C + I + G

C = 480 + 0.5Yd = 480 + 0.5(Y-T) = 480 + 0.5(Y – 70)

Y = 480 + 0.5(Y-70) + 110 + 250

Y = 480 + 0.5Y – 35 + 110 + 250

Y – 0.5Y = 480 + 110 + 250 – 35

0.5Y = 805

Y = 805/0.5 = 1,610

Equilibrium disposable income (Y) = 1,610

 

3.). Equilibrium consumption:

C = 480 + 0.5Yd = 480 + 0.5(Y-T) = 480 + 0.5(Y – 70)

Y = 1,610

C = 480 + 0.5(1610 – 70)

C = 480 + 770 = 1,250

C = 1,250

Equilibrium consumption (C) = 1,250

 

4.). Private savings = Y – T – C

Private savings = 1,610 – 70 – 1,250 = 290

 

Public savings = T – G

Public savings = 70 – 250 = -180

 

Investment spending = Y – C – G

Investment spending = 1,610 – 1,250 – 250 = 110